# The Male CEO and the Female Assistant: Probing Gender Biases in Text-To-Image Models Through Paired Stereotype Test

### v1

@@ Abstract
Recent large-scale Text-To-Image (T2I) models such as DALLE-3 demonstrate great potential in new applications, but also face unprecedented fairness challenges. Prior studies revealed gender biases in single-person image generation, but T2I model applications might require portraying two or more people simultaneously. Potential biases in this setting remain unexplored, leading to fairness-related risks in usage. To study these underlying facets of gender biases in T2I models, we propose a novel Paired Stereotype Test (PST) bias evaluation framework. PST prompts the model to generate two individuals in the same image. They are described with two social identities that are stereotypically associated with the opposite gender. Biases can then be measured by the level of conformation to gender stereotypes in generated images. Using PST, we evaluate DALLE-3 from 2 perspectives: biases in gendered occupation and biases in organizational power. Despite seemingly fair or even anti-stereotype single-person generations, PST still unveils gendered occupational and power associations. Moreover, compared to single-person settings, DALLE-3 generates noticeably more masculine figures under PST for individuals with male-stereotypical identities. PST is therefore effective in revealing underlying gender biases in DALLE-3 that single-person settings cannot capture. Our findings reveal the complicated patterns of gender biases in modern T2I models, further highlighting the critical fairness challenges in multimodal generative systems.
The Male CEO and the Female Assistant: Probing Gender Biases in Text-To-Image Models Through Paired Stereotype Test
Yixin Wan and Kai-Wei Chang University of California, Los Angeles {elaine1wan, kwchang}@cs.ucla.edu
@@ 1 Introduction
Text-To-Image (T2I) models generate images based on textual inputs. Recent development of large T2I models such as Stable Diffusion Rombach et al. (2021), OpenAI’s DALLE-3 OpenAI (2023), and Google’s Imagen Saharia et al. (2022) demonstrate impressive image generation ability, empowering multiple new applications such as AI-assisted content creation [^1] Coca-Cola uses Stable Diffusion to create an advertisement. https://www.youtube.com/watch?v=VGa1imApfdg [^2] Salesforce introduces AI-assisted marketing tools to assist marketers with personalized content creation. https://www.salesforce.com/blog/ai-marketing/. However, these systems suffer from significant fairness challenges. Previous studies have explored gender biases in T2I models Bansal et al. (2022); Bianchi et al. (2023); Naik and Nushi (2023); Friedrich et al. (2023); Cho et al. (2023); Orgad et al. (2023); Wang et al. (2023); Seshadri et al. (2023); Fraser et al. (2023) in single-person generation, in which models are prompted to generate a single individual in images. Several works Bianchi et al. (2023); Orgad et al. (2023); Wang et al. (2023); Friedrich et al. (2023) identified that such biases might be directly propagated from models’ training data, which consists of an enormous amount of internet-scraped single-person images that already carry stereotypes. Since modern T2I models can generate even higher-quality and more complicated images where multiple people might exist, we argue that evaluations on single-person generation are not sufficient for probing intricate patterns of gender biases in T2I systems.
@@
We propose Paired Stereotype Test (PST) as a first step to assess gender biases in T2I models in multi-people generation, which was overlooked by previous works. PST prompts T2I models to generate two individuals in the same image. The two individuals are described with different social identities that are stereotypically associated with the opposite gender. For instance, in the bottom leftmost image in Figure 1, the two individuals are assigned “assistant” and “CEO” as their occupational identities, which are identified by prior work Zhao et al. (2018) to be socially female-stereotyped and male-stereotyped, respectively. Using PST, we conducted extensive experiments to explore two aspects of gender biases in the DALLE-3 model: biases in gendered occupation and biases in organizational power.
Our results show that 1) DALLE-3 is heavily biased in both gender-occupation and gender-power associations, 2) biases in DALLE-3’s generations align with real-world gendered stereotypes, and 3) PST can reveal implicit biases of the T2I model that single-person evaluation settings fail to capture. Interestingly, we observe that DALLE-3 is more likely to portray individuals with male-stereotypical identities as masculine in PST than in single-person setting. Experiment conclusions demonstrate the effectiveness of the proposed PST framework in uncovering the underlying patterns of gender biases in T2I models. Moreover, our findings further highlight the importance and imminence of comprehensively studying and mitigating biases in multimodal generative models.
@@ 2.1 Gender Biases in T2I Models
Previous studies have explored gender biases in T2I models in the single-person setting, where models are prompted to generate images of a single person with specific demographic traits, such as profession or racial group. For instance,  Naik and Nushi (2023) identified the over-representation of white individuals in model-generated images.  Cho et al. (2023) discovered that models might differ on the level and direction of gendered occupation biases. They showed that when generating images for the ‘singer’ profession, minDALL-E Lai et al. (2023) exhibits a propensity towards generating more males, while the Karlo Donghoon Lee and Kim (2022) and Stable Diffusion Rombach et al. (2021) are inclined to generate female images.
However, prior studies’ evaluations are limited to the task setting of generating one individual in the image, whereas contemporary applications of T2I models might involve the generation of multiple people  [^3] For instance, see discussions on multi-character generation in OpenAI Community Dicussion: http://tinyurl.com/4puacskp.. Potential underlying biases in multi-people generations remain unexplored. Additionally, only a few prior researches drew connections between biases demonstrated in model behaviors and in the real world.  Bianchi et al. (2023) is amongst the first to bridge the gap between the study of gender biases in models and in human society. They discovered that the Stable Diffusion (SD) Model Rombach et al. (2021) amplifies biases in gendered occupation from real-world statistics, when generating images of individuals with numerous professions. Inspired by their work, we ground our analysis of gender biases in gendered occupation and in organizational power on identified gender stereotypes in empirical statistics and social science works U.S. Bureau of Labor Statistics (2024); Singh et al. (2019); Kay et al. (2015); Halford and Leonard (2001); Lahtinen and Wilson (1994); Collinson and Hearn (1996).
@@ 3 Method
We propose Paired Stereotype Test (PST) to reveal gender biases in T2I models under challenging generation settings. First, we define the $2$ aspects of gender biases in T2I models that we study. Next, we introduce the proposed PST method and establish $2$ Stereotype Test Score (STS) metrics.
@@ 3.1 Bias Definition
We formally define gender biases in T2I models as the over-generalized association between social identities and specific gender traits in image generation. Such social identities can manifest in a variety of forms, such as professions, racial groups, social status, etc, resulting in multifaceted and intricate perspectives of bias. In our work, we mainly seek to investigate Biases in Gendered Occupation and Biases in Organizational Power. Ideally, a fair model should be equally likely to generate individuals with either gender traits, regardless of their occupation or power levels.
@@ 3.1.1 Biases in Gendered Occupation
We define gender biases in Gendered Occupation as the over-generalized association between gender-stereotypical occupations and gender traits of the generated individual. For instance, consider an image generation task in which a T2I model is given a prompt to generate individuals with the occupational identity of a “hairdresser”. Under this setting, a T2I model is biased if it is more likely to generate an individual with feminine traits than with masculine traits (as illustrated in Figure 2), or tends to associate the “hairdresser” identity with “feminine” gender traits.
@@ 3.1.2 Biases in Organizational Power
Our study is amongst the first to systematically define gender biases in organizational power as the over-generalized association between masculinity and powerful organization roles, as well as between femininity and powerless roles. For example, if a T2I model is disproportionally likely to portray leadership roles like “manager” as individuals with masculine traits than with feminine traits, then it is demonstrating gender biases in organizational power.
3.2 Evaluating Gender Biase in T2I Models
Prior works evaluated gender biases in T2I models under single-person image generation settings Radford et al. (2021); Bianchi et al. (2023); Wang et al. (2023); Naik and Nushi (2023); Cho et al. (2023), in which the models are prompted to portray only one individual with certain social identity. This limitation is partially due to the weaker generation capabilities of prior T2I models- for example, Appendix D demonstrates how Stable Diffusion 2 fails to generate images in T2I setting. However, given the increasingly stronger image generation abilities of T2I systems, applications involving multiple characters are becoming more popular. Therefore, single-person settings are insufficient to capture underlying patterns of gender biases in challenging applications of modern T2I models.
@@ 3.3 Paired Stereotype Test
We propose the Paired Stereotype Test (PST) framework to probe for underlying manifestations of gender biases in T2I systems. PST prompts T2I models to generate two individuals with stereotypical social identities of the opposite genders. For instance, given that managerial organizational positions are stereotypically associated with males and lower-level roles with females, PST will formulate a setting for models to generate one individual of managerial position and another individual of lower-level position in the same image. Intuitively, the paired setting tests if models assume over-generalized gendered associations for stereotypical identities. To quantify gender biases in PST and to validate its effectiveness in revealing underlying gender biases, we report the Stereotype Test Score (STS) as the Absolute level of gender biases of T2I models under the PST setting, on both macroscopic and microscopic levels.
3.3.1 Overall STS
For macroscopic analysis, we report the overall STS averaged on all generated images for all paired stereotypical identities. Let $id_{m}=\{id_{m,1},id_{m,2},...\}$ and $id_{f}=\{id_{f,1},id_{f,2},...\}$ be the $2$ sets of gender-stereotypical identities for males and females. Then, for a T2I model $M$, the generated image when prompted with the paired identities $(id_{m,1},id_{f,1})$ can be represented as $M(id_{m,1},id_{f,1})$. For each of the two individuals in the generated image, we assess if their gender traits conform to the gendered stereotype associated with the assigned social identity. Let $Gender(\cdot,\cdot)$ be a function that takes a generated image and the identity of one individual in the image as inputs, and classifies the gender trait of the individual as $masculine$ or $feminine$. Then, we can define a binary function $Stereo(g_{img},g_{stereo})$ that takes the classified gender trait of the individual in the generated image and the stereotypical gender of the assigned social identity as inputs. The function then determines if the gender trait of the generated individual conforms to the socially stereotyped gender. Specifically,
@@ 3.3.1 Overall STS
The overall STS score can be defined as the average level of gender stereotype conformity for all individuals in generated images for all identity pairs:
$\text{STS}_{ovl}=\frac{1}{4\lvert id_{m}\rvert\cdot\lvert id_{f}\rvert}\sum_{id_{m,i}\in id_{m}}\sum_{id_{f,j}\in id_{f}}$ (1)
@@
3.3.2 Micro STS
Micro STS stratifies bias analysis on the social identity level, and is defined to be the averaged level of stereotype conformity for all individuals with a specific identity in generated images. For example, for a specific male-stereotyped social identity, e.g. $id_{m,1}$, Micro STS is calculated as:
$\text{STS}_{micro}(id_{m,1})=\frac{1}{2\lvert id_{f}\rvert}\sum_{id_{f,j}\in id_{f}}$ (2)
$\;\;\;\;\;\Big{(}Stereo(Gender(M(id_{m,1},id_{f,j}),id_{m,1}),\text{masculine})$
$\;\;\;\;\;+Stereo(Gender(M(id_{f,j},id_{m,1}),id_{m,1}),\text{masculine})\Big{)}$
@@ 4.1 Model Selection
We explore gender biases in OpenAI’s DALLE-3 model and an off-the-shelf Stable Diffusion (SD) model  [^4] https://huggingface.co/stabilityai/stable-diffusion-2. Since the SD model fails to generate most images with two individuals in the paired scenario, we finally choose DALLE-3 for further bias evaluation experiments. Appendix D demonstrates a few failure cases in SD’s generations.
4.2 Evaluation Dataset
Biases in Gendered Occupation To measure biases in gendered occupation, we follow the lists of gender-stereotyped occupations used in  Zhao et al. (2018)’s work, which was originally preprocessed from data provided by the U.S. Bureau of Labor statistics. Complete lists of $20$ male-stereotyped and $20$ female-stereotyped occupations for bias evaluation are shown in Appendix A.1, Table 7.
Biases in Organizational Power To investigate gender biases in organizational power, we combine the aforementioned lists of occupations and several hand-crafted descriptor roles with different power levels in companies. Lists of powerful and powerless descriptors are shown in Appendix A.2, Table 8. To avoid confusion, we take out $4$ occupations that are already indicative of organizational power: “manager”, “supervisor”, “ceo”, and “assistant”. The remaining $36$ occupations are combined with power descriptors in further analyses.
@@ 4.3 Image Generation
Single-Person Setting We first generate images under the single-person setup to establish a direct comparison between gender biases detected using the proposed PST method and the single-person used in prior works. For each of the stereotypical identities (i.e. occupations and corporational roles with different power levels), we prompt DALLE-3 to generate images of an individual with the specified social trait. To ensure the significance of the results, we individually sample $3$ images for the same social identity. The prompt used for generating images is provided in Appendix B, Table 9.
Paired Stereotype Test Under the PST setting for gender bias evaluation, we prompt DALLE-3 to generate images containing two individuals. The two individuals are assigned with stereotypical identities for the opposite gender (i.e. one male-stereotypical identity and one female-stereotypical identity). For biases in gendered occupation, we iterate through all pairwise combinations of male-stereotypical and female-stereotypical professions to get $400$ occupation pairs. To ensure the significance of results, we flip the order of each pair of occupations once to obtain $2$ individual generations, resulting in $800$ generations. For biases in organizational power, we randomly choose $1$ role with high power and $1$ with low power to combine with each occupation, resulting in a total of $36$ pairs of roles with opposite power levels. For each powerful-powerless pair, we flip the order once and individually sample $3$ times to ensure significance, resulting in $216$ generations. Image generation prompts are provided in Appendix B, Table 9.
4.4 Image Annotation
We use Amazon Mechanical Turk Crowston (2012) for hiring human annotators to classify the generated images. For a target individual with an assigned social identity in an image, we hire $3$ annotators to classify the appearance of the individual as demonstrating “masculine” or “feminine” traits. Detailed instructions for the annotators are provided in Appendix C.2. Table 1 demonstrates the overall inter-annotator agreement level under single-person and PST settings for both aspects of gender biases. The high Fleiss’ Kappa score among annotators reflects a satisfactory level of inter-annotator agreement and validates the high quality of labeled data. We also explored the option of using a Visual Question Answering (VQA) model as an automated gender trait classifier in Appendix E. However, the VQA model demonstrates extremely poor performance for images generated in the proposed paired setting, achieving almost 00 agreement level with human-annotated labels.
Bias Aspect Setting Fleiss’ Kappanormal-$\uparrow$
Gendered Occupation
Single
$0.88$
Double
$0.89$
Organizational Power
Single
$0.92$
Double
$0.92$
4.5 Results
4.5.1 Overall STS
We first explore the overall level of gender biases in Occupational Association and Organizational Power, in both single-person and PST settings. Table 2 shows that overall STS scores are higher in the proposed paired setting for both bias aspects. This shows that the proposed PST framework can probe for implicit patterns of gender biases that fail to manifest in the single-person evaluation setting. Below, we further dissect microscopic analyses of the two aspects of bias.
Bias Aspect Stereotype Test Overall STSnormal-$\downarrow$
Gendered Occupation
Single
$10.00$
Paired
$\textbf{47.38}_{\;\;{\color[rgb]{1,0,0}(\textbf{37.38}\uparrow)}}$
Organizational Power
Single
$4.62$
Paired
$\textbf{18.98}_{\;\;{\color[rgb]{1,0,0}(\textbf{14.36}\uparrow)}}$
Table 2: The overall level of gender biases measured by STS. The proposed PST setting is able to reflect a higher level of underlying biases in DALLE-3, compared to the single-person evaluation setting.

### v3

# The Male CEO and the Female Assistant: Evaluation and Mitigation of Gender Biases in Text-To-Image Generation of Dual Subjects

@@ Abstract
Recent large-scale T2I models like DALLE-3 have made progress in reducing gender stereotypes when generating single-person images. However, significant biases remain when generating images with more than one person. To systematically evaluate this, we propose the Paired Stereotype Test (PST) framework, which queries T2I models to depict two individuals assigned with male-stereotyped and female-stereotyped social identities, respectively (e.g. “a CEO” and “an Assistant”). This contrastive setting often triggers T2I models to generate gender-stereotyped images. Using PST, we evaluate two aspects of gender biases – the well-known bias in gendered occupation and a novel aspect: bias in organizational power. Experiments show that over 74% images generated by DALLE-3 display gender-occupational biases. Additionally, compared to single-person settings, DALLE-3 is more likely to perpetuate male-associated stereotypes under PST. We further propose FairCritic, a novel and interpretable framework that leverages an LLM-based critic model to i) detect bias in generated images, and ii) adaptively provide feedback to T2I models for improving fairness. FairCritic achieves near-perfect fairness on PST, overcoming the limitations of previous prompt-based intervention approaches.
The Male CEO and the Female Assistant: Evaluation and Mitigation of Gender Biases in Text-To-Image Generation of Dual Subjects
Yixin Wan  and Kai-Wei Chang Department of Computer Science University of California, Los Angeles {elaine1wan, kwchang}@cs.ucla.edu
@@ 1 Introduction
Text-To-Image (T2I) models generate images based on textual prompts. T2I models like OpenAI’s DALLE-3 OpenAI (2023) demonstrate impressive image generation ability, empowering multiple new applications such as AI-assisted content creation [^1] Coca-Cola uses Stable Diffusion as a novel tool to create an advertisement. [^2] Salesforce introduces AI-assisted marketing tools to assist marketers with personalized content creation.. However, several studies have revealed that when prompted to generate images of a person with a specific occupation (e.g., doctor or nurse), T2I models often exhibit gender biases  Bansal et al. (2022); Bianchi et al. (2023); Naik and Nushi (2023); Friedrich et al. (2023); Cho et al. (2023); Orgad et al. (2023); Wang et al. (2023); Seshadri et al. (2023); Fraser et al. (2023). With the development of various mitigation methods (e.g.,  (Bansal et al., 2022)), recent T2I models like DALLE-3 have made notable improvements in generating more diverse images. For instance, the top row in Figure 1 shows how DALLE-3 outputs anti-stereotypical genders in single-person images.
However, many applications of T2I models—such as AI-generated visual advertisements, movie frames, game designs—require the generation of images with multiple people. Therefore, it is crucial to evaluate the intricate gender biases of T2I models in multi-person scenarios, as the relationships between individuals can influence model behavior and make debiasing even more challenging.
@@
Figure 2: Visualization of the evaluation framework used in PST.
In this study, we address this research gap and propose the Paired Stereotype Test (PST) to evaluate gender biases in T2I models in a challenging dual-subject (two people) image generation. PST prompts T2I models to generate one individual with male-stereotypical attribute and another with female-stereotypical attribute. For instance, in the bottom leftmost image in Figure 1, the two individuals are assigned “assistant” and “CEO” as their occupational identities, which are identified by prior work Zhao et al. (2018) to be socially female-stereotyped and male-stereotyped, respectively. As we can observe from the examples in the bottom row of Figure 1, the setting in PST is especially challenging for T2I models, as models easily follow both social stereotypes in this contrastive scenario: DALLE-3 default to depicting a male CEO and a female assistant, a male manager and a female secretary, etc.
To support the systematical analysis of T2I models using PST, we collect and contribute 1952 prompts to evaluate two types of gender biases: bias in occupational association, and a novel aspect of bias in organizational power, which studies the unfair correlation of males with high-power organizational roles (e.g. CEO), and females with low-power roles (e.g. assistant). Additionally, we propose the Stereotype Score (SS) to quantify gender bias via levels of stereotype adherence.
Using PST, we conduct extensive experiments to unveil gender bias in OpenAI’s DALLE-3 model. Our results show that 1) DALLE-3 is heavily biased in both gender-occupation and gender-power associations, 2) biases in DALLE-3’s generations align with real-world gendered stereotypes, and 3) PST can reveal implicit biases of the T2I model that single-subject evaluation settings fail to capture. Interestingly, we observe that DALLE-3 is more likely to propagate male stereotypes in PST than in single-subject settings.
Furthermore, we explored approaches to resolve the newly observed biases in PST setting. Contrary to public beliefs, previously widely adopted prompt-based mitigation methods could result in overshooting biases, creating the opposite stereotype. This warns that vanilla prompt engineering technique fails to address biases in complex T2I tasks controllably. To address this drawback, we propose a novel framework called FairCritic, which adopts an LLM-based critic model to judge the fairness of generated images and adaptively provide concrete feedback to augment the generation context and mitigate observed biases. Results show that FairCritic achieves near-perfect fairness outcomes under PST, remarkably surpassing vanilla prompt-based interventions. What’s more, FairCritic utilizes a fully interpretable framework, allowing for easy analysis of intermediate fairness feedback provided by the critic model. The method is also promising in generalizing to even more challenging image generation tasks.
Our PST framework succeeds in uncovering concealed gender biases in T2I models that fail to manifest in single-subject generation settings, and our novel FairCritic approach effectively and interpretably addresses the bias problem in complicated generation settings. We will release our code and data upon acceptance.
@@ 2.1 Gender Biases in T2I Models
Previous studies have explored gender biases in T2I models in the single-subject generation setting, where models are prompted to generate images of a single person with specific demographic traits, such as profession or racial group. For instance,  Naik and Nushi (2023) identified the over-representation of white individuals in model-generated images.  Cho et al. (2023) discovered that models might differ on the level and direction of gendered occupation biases. They showed that when generating images for the ‘singer’ profession, minDALL-E Lai et al. (2023) exhibits a propensity towards generating more males, while the Karlo Lee et al. (2022) and Stable Diffusion Rombach et al. (2021) are inclined to generate female images. Several works Bianchi et al. (2023); Orgad et al. (2023); Wang et al. (2023); Friedrich et al. (2023) identified that such biases might be directly propagated from models’ training data, which consists of an enormous amount of internet-scraped single-person images that already carry stereotypes.
However, prior studies’ evaluations are limited to the task setting of generating one individual in the image, whereas contemporary applications of T2I models might involve the generation of multiple people  [^3] For instance, see discussions on multi-character generation in OpenAI Community.. Underlying biases in dual-subject or even multi-subject generations remain unexplored. Additionally, only a few prior researches drew connections between biases demonstrated in model behaviors and in the real world.  Bianchi et al. (2023) is amongst the first to bridge the gap between the study of gender biases in models and in human society. They discovered that the Stable Diffusion (SD) Model Rombach et al. (2021) amplifies biases in gendered occupation from real-world statistics, when generating images of individuals with numerous professions. Inspired by their work, we ground our analysis of gender biases in gendered occupation and in organizational power on identified gender stereotypes in empirical statistics and social science works U.S. Bureau of Labor Statistics (2024); Singh et al. (2019); Kay et al. (2015); Halford and Leonard (2001); Lahtinen and Wilson (1994); Collinson and Hearn (1996).
@@ 3 Method
We propose Paired Stereotype Test (PST) to reveal gender biases in T2I models under challenging generation settings. First, we define the $2$ aspects of gender biases in T2I models that we study. Next, we introduce the proposed PST method and establish $2$ Stereotype Score (SS) metrics.
@@ 3.1 Bias Definition
We formally define gender biases in T2I models as the over-generalized association between social attributes and specific gender traits in image generation. Such social attributes can manifest in a variety of forms, such as professions, social status, etc, resulting in multifaceted and intricate biases. We mainly seek to investigate Biases in Gendered Occupation and Biases in Organizational Power. Ideally, a fair model should be equally likely to generate individuals with either gender traits, regardless of their occupation or power levels.
@@ 3.1.1 Biases in Gendered Occupation
Consider an image generation task in which a T2I model is given a prompt to generate individuals with the occupational identity of a “hairdresser”. Under this setting, a T2I model is biased if it is more likely to generate an individual with feminine traits than with masculine traits (as illustrated in Figure 3), or tends to associate the “hairdresser” identity with “feminine” gender traits.
@@ 3.1.2 Biases in Organizational Power
Our study is amongst the first to systematically study gender bias in organizational power. We define this bias as the over-generalized association of masculinity with high-power organization roles, and femininity with low-power roles. A T2I model is biased in the organizational power aspect if, within the same industry (e.g. accounting), it is disproportionally likely to portray leadership roles (e.g. accounting manager) as individuals with masculine traits and low-power roles (e.g. accounting assistant) as with feminine traits.
@@ 3.2 Paired Stereotype Test
We propose the Paired Stereotype Test (PST) framework to probe for gender biases in dual-subject T2I generation.
3.2.1 Task Formulation
PST prompts T2I models to simultaneously generate one individual with male-stereotypical social attributes and another with female-stereotypical attributes. For instance, given that the occupation “CEO” is stereotypically associated with males and “assistant” with females, PST formulates a setting to prompt models to generate a CEO and an assistant in the same image. Intuitively, this contrastive dual-subject-based setting is challenging for T2I models—models easily adhere to both stereotypes for the opposite genders, e.g. generating a male CEO and a female assistant.
3.2.2 Evaluation Prompts
For evaluating the 2 aspects of gender bias under PST, we designed 944 template-based prompts using social attribute descriptors stereotypically associated with different genders. Details on prompt construction are provided in Appendix A.2.1.
3.2.3 Evaluation Metrics
To quantify gender bias in generated images under the PST setting, we propose and report the Stereotype Score (SS) as the Absolute level of gender stereotype adherence in model generations.
Figure 4: Gender Ratio of Biased Images.
Figure 5: Bias Amplification and Gender Ratio of Biased Images.
Stereotype Score (SS)   We define SS as the ratio gap between stereotype-adhering individuals and anti-stereotype individuals in model-generated images. For instance, the overall SS for occupational gender bias can be calculated as:
Let $id_{m}=\{id_{m,1},id_{m,2},...\}$ and $id_{f}=\{id_{f,1},id_{f,2},...\}$ be $2$ sets of gender-stereotypical occupations for males and females, respectively. Let $M$ be a T2I model. Then, when prompted to create an image with two individuals with occupation attributes $(id_{m,1},id_{f,1})$, model generation can be represented as $M(id_{m,1},id_{f,1})$. Let $Gender(\cdot,\cdot)$ be a function that takes a generated image and the position of one individual in the image as inputs, and classifies the gender trait of the individual as $masculine$ or $feminine$. Then, we can define a binary function $Stereo(g_{img},g_{stereo})$ that determines whether the classified gender trait of an individual with a specific occupation adheres to the socially stereotyped gender of the occupation. Specifically,
@@ 3.2.3 Evaluation Metrics
The overall SS score for occupational bias can then be calculated as the average level of gender stereotype adherence across all generated individuals:
$\text{SS}_{ovl}=\frac{1}{4\lvert id_{m}\rvert\cdot\lvert id_{f}\rvert}\sum_{id_{m,i}\in id_{m}}\sum_{id_{f,j}\in id_{f}}$ (1)
@@
Additionally, we also calculate the SS for stratified aspects, i.e. for male-stereotyped and female-stereotyped attributes, separately.
@@ 4.1 Model Selection
We explore gender biases in OpenAI’s DALLE-3 model and an off-the-shelf Stable Diffusion (SD) model  [^4] https://huggingface.co/stabilityai/stable-diffusion-2 with default hyperparameter settings. Since the SD model fails to generate most images with two individuals in the paired scenario, we finally choose DALLE-3 for further bias evaluation experiments. Appendix C demonstrates a few failure cases in SD’s generations.
@@ 4.2 Image Generation
Single-Subject Generation We establish the single-subject generation setting explored in previous works Radford et al. (2021); Bianchi et al. (2023); Wang et al. (2023); Naik and Nushi (2023); Cho et al. (2023), in which models are asked to depict only one individual in images, as the baseline setting for comparison. We prompt DALLE-3 to generate single-person images for stereotypical attributes like occupations and organizational roles. To ensure the significance of results, we individually sample $3$ images for the same social identity.
Paired Stereotype Test We follow the task description for PST in Section 3.2.1 to generate images for evaluation. For biases in organizational power, due to computational limits, we conducted $3$ independent runs, during each we sampled 1 pair of high-power level and low-power level roles for each occupation. This results in a total of $216$ generations.
4.3 Gender Annotation
To annotate the gender of individuals in generated images, we mainly use human annotations from the Amazon Mechanical Turk Crowston (2012) platform. Annotation details and statistics are provided in Appendix B. Additionally, we explored 2 automated options for gender annotation: (1) the Bootstrapping Language-Image Pre-training (BLIP) Visual Question Answering (VQA) model Li et al. (2022), and (2) using the FairFace (Kärkkäinen and Joo, 2019) classifier. The VQA model demonstrates extremely poor performance for images generated in the proposed paired setting, achieving almost 00 agreement level with human-annotated labels. FairFace classifier achieves good alignment with human-annotated results, and is promising as an automated annotation tool in future studies. Performance details are in Appendix D.
4.4 Experiment Results
4.4.1 Overall Bias
We first explore the overall level of gender biases in Occupational Association and Organizational Power. Table 1 shows that overall SS scores, which are averaged across all generations for each bias aspect, are higher in PST than in the single-subject generation setting. This shows that PST unveils concealed patterns of gender biases that fail to manifest in evaluation methods under single-person generation settings. We visualize the percentage of biased images across model generations in Figure 6—PST(Both) identifies cases where both individuals in generated images adhere to gender stereotypes, whereas PST (Any) identifies cases where either of the generated figures are stereotyped. We observe a notable increase in the ratio of stereotyped or biased generations under PST.
Figure 6: Percentage of biased images across different task settings and aspects of biases.
Figure 7: Occupation-level SS under the PST setting by occupations, sorted in ascending order.
Figure 8: PST-revealed biases align with real-world gender segregation across occupations. Each pair of bars in the figure represents the percentage of feminine figures in generated images and the ratio of females in real-world labor statistics for a specific profession.